### v2

# Cascades towards noise-induced transitions on networks revealed using information flows

Abstract
[…]
00footnotetext: Corresponding author: c.vanelteren@uva.nl11footnotetext: Computational Science Lab, University of Amsterdam, The Netherlands22footnotetext: Institute for Advanced Study, University of Amsterdam, Amsterdam, The Netherlands33footnotetext: Polder center, Institute for Advanced Study, The Netherlands44footnotetext: Complexity Science Hub Vienna, Vienna, Austria
1 Introduction
Multistability is a crucial characteristic observed in various real-world complex systems [Ladyman2013, vanNes2016]. It describes the phenomenon in which a system, under the influence of noise, explores different states within its state space at various spatio-temporal scales. Examples range from the brain’s ability to operate in one cognitive mode for an extended period before rapidly changing its state [Kandel2000, Fries2015]. Larger-scale instances include multistability in opinion dynamics [Galam2020] and ecosystems or climate systems balancing between metastable states [1986, Wunderling2021].
Evidence increasingly suggests that noise fluctuations play a fundamental role in transitions between attractor states [Beggs2012, Mitchell1993, Mitchella, Forgoston2018a]. Traditionally, noise is perceived as an unwanted signal, diminishing system effectiveness or signal quality. However, in nonlinear systems, such as interconnected neurons in the brain, noise can play a surprising role by enabling the exploration of larger parts of the state space, allowing the system to escape local minima [Czaplicka2013, Nicolis2016].
Here, we focus on a particular subset of dynamical systems where each node’s state in the next time step is governed by the Boltzmann-Gibbs distribution, akin to a local thermal equilibrium perspective. Such models have been applied to describe various behaviors, including neural dynamics [Hopfield1982b], opinion dynamics, ferromagnetic spins [Glauber1963], and organized criminal gang interactions [DOrsogna2015a].
Dynamical networks governed by the Boltzmann-Gibbs distribution involve nodes choosing a new state based on the local equilibrium potential induced by their neighbors. The transition probability from state $s_{i}$ to $s_{i}^{\prime}$ is given by:
$P(s_{i}\to s_{i}^{\prime})=\frac{1}{1+\exp(-\beta\Delta E(s_{i},s_{i}^{\prime}))},$ (1)
where $\Delta E(s_{i},s_{i}^{\prime})$ represents the energy difference for the state transition, and $\beta$ is the noise parameter.
In physical applications, this transition involves overcoming an energy barrier (fig. 1), akin to frustration, assortativity, or a fitness score. A global ‘temperature’ parameter introduces noise, influencing the likelihood of choosing higher energy states. For networked dynamical system, the structure itself significantly influences systemic transitions, making them more likely than in systems without network effects [Harush2017a, Gao2016, Wunderling2020, Wunderling2021]. Our method uncovers the network percolation process facilitating endogenously generated, noise-induced transitions, offering a different perspective from traditional approaches. This computational method requires only access to system state probabilities over time, making it broadly applicable.
For systems characterized by low noise levels and governed by the Boltzmann-Gibbs distribution, metastable behavior arises due to the presence of multiple local minima (see fig. 1). This study focuses on elucidating two crucial roles pivotal in facilitating distinct phases of metastable transitions within networked systems.
Firstly, we introduce the concept of initiator nodes, which play a fundamental role in propagating noise within the system. These initiator nodes serve to destabilize nodes occupying lower energy states, thereby priming the system towards a transition. As the transition initiates, a cascade of flips is rapidly triggered by these initiator nodes, propelling the system towards a new metastable state. Secondly, we identify the significance of stabilizing nodes in counteracting the effects of noise and maintaining the system in a metastable state. These nodes act to stabilize the system amidst the disruptive influence of noise, thereby contributing to the persistence of the metastable state.
To quantitatively assess the impact of correlations at both short and long time scales, two novel information theoretic measures are introduced. These measures, namely integrated mutual information and asymptotic information, are designed to capture the effects of short-term and long-term correlations, respectively. They provide valuable insights into the dynamics of metastable transitions within networked systems and can be directly computed on observations.
Figure 1: A dynamical network governed by kinetic Ising dynamics produces multistable behavior. (a) A typical trajectory is shown for a kite network for which each node is governed by the Ising dynamics with $\beta\approx 0.534$. The panels show system configurations $S_{i}\in S$ as the system approaches the tipping point (orange to purple to red). For the system to transition between attractor states, it has to cross an energy barrier (c). (b) The dynamics of the system can be represented as a graph. Each node represents a system configuration $S_{i}\in S$ such as depicted in (a). The probability for a particular system configuration $p(S)$ is indicated with a color; some states are more likely than others. The trajectory from (a) is visualized. Dynamics that move towards the tipping point (midline) destabilize the system, whereas moving away from the tipping point are stabilizing dynamics. (c) The stationary distribution of the system is bistable. Crossing the tipping point requires crossing a high energy states (dashed line). Transitions between the attractor states are infrequent and rare. For more information on the numerical simulations see A.2.
2 Methods
Fluctuations and their correlations at time $\tau$ are captured using Shannon’s mutual information [Cover2005] shared between a node’s state ($s_{i}^{t}$) at time $t$ and the entire future system state ($S^{t+\tau}$). $I(s_{i}^{\tau}:S^{\tau+t})$. The time lag $t$ is used to analyze two key features of information flows of a system: the area under the curve (AUC) of short-term information, and sustained level of long term information.
The contribution of a node to the dynamics of the system will differ depending on the network connectivity of a node (fig. A.8) [vanElteren2022, Quax2013]. The total amount of fluctuations shared between the node’s current state and the system’s short-term future trajectory is computed as the integrated mutual information
$\begin{split}\mu(s_{i})=\sum_{t=0}^{\infty}(I(s_{i}^{\tau}:S^{\tau+t})-\omega_{s_{i}})\Delta t.\end{split}$ (2)
Intuitively, $\mu(s_{i})$ represents a combination of the intensity and duration of the short-term fluctuations on the (transient) system dynamics [vanElteren2022]. It reflects how much of the node state is in the “working memory” of the system.
The term $\omega(s_{i})\in\mathbb{R}_{\geq 0}$ quantifies the long term memory of the system. Positive values of $\omega(s_{i})$ imply a separation of time-scales: short-lived correlations dissipate for increasing $t$ and are replaced by slower system fluctuations. These slow fluctuations embody the multiple attractor states the system can inhabit. Around a low energy state, the system produces short-lived fluctuations. As the system approaches a tipping point a new low energy state will be chosen. Correspondingly, the correlations of a node with this new future system state will produce long(er) timescale correlations. The next tipping point will be reached on a much longer timescale.
Using these information features, each node can be assigned to a different role based on their contribution to the metastable transition. We denote nodes with short-lived correlations as initiators pushing nodes towards a tipping point. In contrast, nodes with longer-lived correlations are referred to as stabilizers. For these nodes, their dynamics are less affected by short-lived correlations, and they require a higher mixing state for to transition from one state to another. The role assignment will be further discussed in 3.1.
The information flows are computed using exact computation. The process involved randomly generating a connected graph of $n=10$ nodes and grouping the states as distance to the tipping point, after which the conditional distribution was evaluated up to $\tau=300$. The tipping point is defined as the energy barrier between two locally stable states. For the Ising model this is the collection of states such that $\langle S\rangle=0.5$. For further details on replicating please see appendix A.
3 Results
In fig. 2a-e the information flows are shown at different stages as the system approaches a tipping point. The metastable transition was decomposed by considering the local information flows from a given system partition $S_{\gamma}=\{S^{\prime}\subseteq S|\langle S^{\prime}\rangle=\gamma\}$ where $\gamma\in[0,1]$ is the fraction of nodes having state 1. This yields the conditional integrated mutual information as
$\begin{split}\mu(s_{i}|\langle S\rangle)=\sum_{t=0}^{\infty}(I(s_{i}^{\tau}:S^{\tau+t}|\langle S^{\tau}\rangle)-\omega_{s_{i}})\Delta t.\end{split}$ (3)
Details about the estimation procedure can be found in appendix: A.5.
Figure 2: (a-e) Information flows as distance to tipping point. Far away from the tipping point most information processing occurs in low degree nodes (f,g). As the system moves towards the tipping point, the information flows increase and the information flows move towards higher degrees. (f) Integrated mutual information as function of distance to tipping point. The graphical inset plots show how noise in introduced far away from the tipping point in the tail of the kite graph. As the system approaches the tipping point, the local information dynamics move from the tail to the core of the kite. (g) A rise in asymptotic information indicates the system is close to a tipping point. At the tipping point, the decay maximizes as trajectories stabilize into one of the two attractor states.
Two things are observed from fig. 2. First, the tipping point is reached by a domino effect where low degree nodes play an “initiator” role early in the process. In this model, low degree nodes are most susceptible to noise fig. A.8 and therefore are more likely to pass on fluctuations to neighbors. Far away from the tipping point (fig. 2a), nodes with lower degree have higher shared information (higher $\mu(s_{i}|\langle S\rangle)$) than higher degree nodes. Lower degree nodes can initiate a metastable transition by injecting noise into the system. Without this injected noise, it would be less likely for a metastable transition to occur. In other words, in a system that is slightly destabilized by low degree nodes with high energy (fluctuating states), the transition towards a tipping point becomes more likely as neighboring higher degree nodes are more likely to become “initiator” nodes. This cascade progresses whereby new initiator nodes are formed through local fluctuations.
Second, an increase in asymptotic behavior correlates with the system transitioning from one attractor state to another. The asymptotic information remains low far away from the tipping point, and monotonically increases as the system approaches the tipping point fig. 2{b, c}). A node’s asymptotic information encodes how much predictive information a node has about the future system state. After a tipping point, the system either relaxes to the closest attractor state or transitions across the tipping point into the next attractor state. After such a transition, the dynamics of the nodes slow down. That is, all but the nodes with the lowest degrees are locally frozen as the system dynamics restabilizes after a noise-induced perturbation. A node with high asymptotic information will have more information regarding which side of the tipping point the system ends up being.
To illustrate what is encoded in the information flows trajectories were computed from the attractor state $S=\{0,\dots,0\}$ and simulated for $t=5$ steps. In fig. 3 a trajectory is shown that maximizes
$\log p(S^{t+1}|S^{t},S^{0}=\{0,\dots,0\},\langle S^{5}\rangle=0.5).$
These trajectories reveal how the information flows measured in fig. 2c are caused by the sequence of flips generated from the “tail” in the kite graph. These tail nodes are uniquely positioned due to their higher potential to pass on fluctuations to their neighbors eventually causing a cascade of flips that reach the tipping point.
The domino effect is not completely correlated with degree. As the system approaches the tipping point, destabilizing fluctuations tend to be caused by lower degree nodes, but as the system approaches the tipping point network effects play a profound role. For example, consider node 8 and node 3. Node 8 has degree 2 and has the highest integrated mutual information when 2 bits are flipped in the system (fig. 2b). The dynamics for node 8 for all states where $\langle S\rangle=0.2$ (or 0.8 by symmetry) indicate that 8 is essential in propagating the fluctuations generated by 9. At the tipping point, node 8 shares the highest information with the system. In contrast, node 3 which has degree 6 has low shared information prior to the tipping, indicating that 3 is less involved with initializing the tipping point. At the tipping point, however, node 3 has high amounts of shared information with the future system states, similar to that of node 8. This makes it hard to generate a strict rule based on network connectivity alone what role a node has in the tipping behavior. Both the network structure and the dynamics fundamentally interact in generating the tipping points. Furthermore, the role of a node may change as the system approaches a tipping point.
Figure 3: The tipping point is initiated from the bottom up. Each node is colored according to state 0 (black) and state 1 (yellow) Shown is a trajectory towards the tipping point that maximizes $\sum_{{t=1}}^{{5}}\log p(S^{{t+1}}|S^{t},S^{0}=\{0\},\langle S^{5}\rangle)=0.5)$. As the system approaches the tipping point, low degree nodes flip first, and recruit “higher” degree nodes to further destabilize the system and push it towards a tipping point. In total 30240 trajectories that reach the tipping point in 5 steps, and there are 10 trajectories that have the same maximized values as the trajectory shown in this figure.
The information flows reflect how certain a given node is about the future system state. At the tipping point the system is most likely to either (a) move from one attractor state to another, or (b) relax back to the attractor state it evolved from (fig. 3). Path analysis revealed that the most likely paths reaching the tipping point from one of the ground state results in a configuration in which a high degree cluster set of nodes has to flip (e.g. 1,0,3,4,6 in fig. 3 at $\langle S\rangle=0.5)$. This trajectory is less likely than essentially reversing the path shown in fig. 3. Hence, most of the tipping points “fail” and relax back to the attractor state from which it evolved (fig. 4b).
Figure 4: (a) Shown are the conditional probability at time $t=10$ relative to the tipping point. The shared information between the hub node 3 and the tail node 8 is shared is similar but importantly caused through different sources. The hub (node 3) has high certainty on that the system macrostate will be the same sign as its state. In contrast, node 8 has high certainty that the system macrostate will be opposite to its state at the tipping point. This is caused by the interaction between the network structure and the system dynamics whereby the most likely trajectories to the tipping point from the stable regime is mediated by the noise-induced dynamics from the tail to the core in the kite graph (see main text).(b) Successful metastable transitions are affected by network structure. Successful metastable transitions are those for which the sign of the macrostate is not the same prior and after the tipping point, e.g. the system going from the 0 macrostate side to the +1 macrostate side or vice versa. Shown here are the number of successful metastable transitions for fig. 5 under control and pinning interventions on the nodes in the kite graph.
The increased information of node 8 around the tipping point can now be understood by considering what kind of information 8 has about the future of the system. The path analysis revealed that the network structure plays a fundamental role whereby a domino effect from the “bottom-up” is the most likely path to and from a tipping point. This implies that the information that node 8 and node 3 store about the future of the system differs but ends up providing the same amount of shared information. In fig. 4a the conditional probabilities are shown of each node relative to the tipping point. Both node 3 and node 8 have the lowest uncertainty about the future system state. However, the nature of this uncertainty differs. Relative to the tipping point, the node 3 has more certainty that the average of the system state will be equal to its state at the tipping point. This reflects the node’s ability to “choose” the next stable point. This is most likely caused for the kite graph by a failure of the system to transition between attractor states (fig. 4b): most transitions are more likely to transition back to the attractor state it evolved from than to another attractor state. Node 8, however, shares different information about the future system state. Figure 4 shows that node 8 has higher certainty that the future system state will most likely have opposite sign to its state at the tipping point. As most tipping points fail to transition between metastable points, node 8 will have the opposite state to what it was at the tipping point. This gives node 8 a non-intuitive high predictive power of the system’s future.
The information flows reflect the most probable trajectories around the partition $\langle S\rangle=c$ and give unique insights into the mechanism driving the tipping behavior. Over time, local clusters will stabilize. Some nodes will experience more “frustration” than others. In other words, the node will tend to change state more as the effect of a node flip percolates through the system. For example, nodes 5 (yellow) and 6 (orange) have the lowest asymptotic information while still having a relatively high degree. These nodes experience more frustration as they attempt to reconcile with the states of the nearest neighbors.
Figure 5: For a system to cross a tipping point two different types of nodes are identified. Stabilizer nodes are essential for system to move from one attractor state to another, they contain information about how the system chooses its next attractor state. Initiator nodes are essential to propagate noise into the system. (a) The effect of pinning intervention per node on 0 state in Erdos-Renyi graphs (N = 100 graphs with 10 nodes each and p = 0.2, 6 different seeds). Shown are normalized system fluctuations (second moment) and time spent below the tipping relative to the control per network. Pinning intervention on initiator nodes increases the occurrence of tipping points. In contrast, interventions on stabilizers prevents tipping points and increases noise above the tipping point. For more details on the role approximation please see 3.1. In (b) typical system trajectories are shown under pinning intervention on a node for the kite graph. Colors per system trajectory reflect intervention on the corresponding color in the kite graph (inset b). Intervention on initiator nodes (e.g. purple / red purple) removes system fluctuations below the tipping point ($\langle S\rangle<0.5$) congruent with the pinned state (0), but increases the system fluctuations above the tipping point compared to control (blue, bottom plot). In contrast, stabilizer nodes (e.g. green) increase system fluctuations compared to control but stabilizes tipping points.
3.1 Role division in tipping behavior
The role for node $i$ can be approximated by the difference between integrated mutual information and asymptotic information
$r_{i}=\max_{\langle S\rangle}\mu^{*}(s_{i}|\langle S\rangle)-\max_{\langle S\rangle}\omega^{*}(s_{i})\in[-1,1],$ (4)
with
$\mu^{*}(s_{i}|\langle S\rangle)$ $=\frac{\mu(s_{i}|\langle S\rangle)}{\max_{j,\langle S\rangle}\mu(s_{j}|\langle S\rangle)}$ (5)
$\omega^{*}(s_{i}|\langle S\rangle)$ $=\frac{\omega(s_{i}|\langle S\rangle)}{\max_{j,\langle S\rangle}\omega(s_{j}|\langle S\rangle)}.$
For role values close to 1, the node is classified as a (pure) initiator. These nodes have high integrated mutual information indicating high predictive information about short-lived system trajectories. However, these nodes lack long-term predictive information about future system states. Conversely, a node classified as -1, has is a pure stabilizer. Roles having a value $r_{i}\sim 0$ are more difficult to interpret as the zero value could be caused by an equally large integrated mutual information and asymptotic information or a generally lacking high score in both.
Simulated interventions were used to validate the role description identified using information features. In fig. 5 the effect on system fluctuations and number of tipping points using simulated interventions were shown. Nodes were pinning to the 0 state in separate simulations in different Erdos-Renyi graphs (N = 100 graphs with 10 nodes, non-isomorphic connected graphs, please see appendix: A.10). From the results two conclusions can be made.
First, tipping points are promoted proportional to a node’s role: the higher a node scores as a initiator, the more likely and intervention will induce tipping points. A typical system trajectory under intervention is shown in fig. 5b. We denote $\langle S\rangle<0.5$ as below the tipping point as interventions were performed by pinning the node state at 0. This would create a bias towards the attractor state $S=\{0,\dots,0\}$. Interventions on initiator nodes mitigate system fluctuations below the tipping point compared (circle markers) to control conditions (blue lines in dashed lines in fig. 5a, and blue line in fig. 5b). In contrast, it increases the system fluctuations above the tipping points (square markers). Together this implies that as a node’s role approaches 1, the role of this node is promoting towards initiating the tipping behavior, but not preventing tipping points from occurring.
Second, interventions on stabilizer nodes are essential to stabilize the transition from one attractor state to the other. Intervening on stabilizer nodes increases the system fluctuations above the tipping point (square markers in fig. 5) and increases the duration below the tipping point. The increased fluctuations indicate that the stabilizer nodes are necessary for the system to successfully transition between attractor states.
4 Discussion
Understanding how metastable transitions occur may help in understanding how, for example, a pandemic occurs, or a system undergoes critical failure. In this paper, dynamical networks governed by the Boltzmann-Gibbs distribution were used to study how endogenously generated metastable transitions occur. The external noise parameter (temperature) was fixed such that the statistical complexity of the system behavior was maximized (see appendix A.2).
The results show that in the network two distinct node types could be identified: initiator and stabilizer nodes. Initiator nodes are essential early in the metastable transition. Due to their high degree of freedom, these nodes are more effected by external noise. They are instigators and inject propagate in the system, destabilizing more stable nodes. In contrast, stabilizer nodes, have low degree of freedom and require more energy to change state. These nodes are essential for the metastable behavior as they stabilize the system macrostate. During the metastable transition a domino sequence of node state changes are propagated in an ordered sequence towards the tipping point.
This domino effect was revealed through two information features unvealing an information cascade underpinning the trajectories towards the tipping point.
Integrated mutual information captured how short-lived correlations are passed on from the initator nodes. In the stable regime (close to the ground state) low degree nodes drive the system dynamics. Low degree nodes destabilize the system, pushing the system closer to the tipping point. In most cases, the initiator nodes will fail in propagating the noise to their neighbors. On rare occasions, however, the cascade is propagated progressively from low degree, to higher and higher degree. A similar domino mechanism was recently found in climate science [Wunderling2020, Wunderling2021]. Wunderling and colleagues provided a simplified model of the climate system, analyzing how various components contribute to the stability of the climate. They found that interactions generally stabilize the system dynamics. If, however, a metastable transitions was initialized, noise was propagated through a similar mechanism as found here. That is, an “initializer” node propagated noise through the system which created a domino effect that percolated through the system. The results from this study mirrors these conclusions and provides a model-free language to express these domino effects.
An increase in asymptotic information forms an indicator of how close the system is to a tipping point. Close to the ground state, the asymptotic information is low, reflecting how transient noise perturbations are not amplified and the system macrostate relaxes back to the ground state. As the system approaches the tipping point, the asymptotic information increases. As the distance to the ground state increases, the system is more likely to transition between metastable states. After the transition, there remains a longer term correlation. Asymptotic information reflects the long(er) timescale dynamics of the system. This “rest” information peaks at the tipping point, as the system chooses its next state.
The information viewpoint uniquely reveals a complex mechanism of interaction underlying the system macrostate. It allows for compressing the high dimension probability distribution in a away to understand what elements are fundamental for a tipping point ot be reached. It revealed how some nodes may have high predictive information, which is hard to infer from their interaction structure alone fig. 4. Integrated information and asymptotic information jointly readout the separation of fast-time scale dynamics that tend to stabilize noise-induced dynamics, and slow timescale dynamics indicating a metastable transition. Importantly, these measures can be directly computed on data.
5 Conclusions
Our information theoretic approach offers an alternative view to understand how metastable transitions are generated by dynamical networks. Two information features were introduced that decompose the metastable transition in sources of high information processing (integrated mutual information) and distance of the system to the tipping point (asymptotic information). A domino effect was revealed, whereby low degree nodes initiate the tipping point, making it more likely for higher degree nodes to tip. On the tipping point, long-term correlations stabilizes the system inside the new metastable state. Importantly, the information perspective allows for estimating integrated mutual information directly from data without knowing the mechanisms that drive the tipping behavior. The results highlight how short-lived correlations are essential to initiate the information cascade for crossing a tipping point.
6 Limitations
Integrated mutual information was computed based on exact information flows. This means that for binary systems it requires to compute a transfer matrix on the order of $2^{|S|}\times 2^{|S|}$. This reduced the present analysis to smaller graphs. It would be possible to use Monte-Carlo methods to estimate the information flows. However, $I(s_{i}^{\tau}:S^{\tau+t})$ remains expensive to compute. When using computational models, it requires to compute the conditional and marginal distributions which are on order $\mathbb{O}(2^{|S|})$ and $\mathbb{O}(2^{t|S|})$ respectively.
In addition, the decomposition of the metastable transition depends on the partition of the state space. Information flows are in essence statistical dependencies among random variables. Here, the effect of how the tipping point was reached was studied by partition the average system state in terms of number of bits flipped. This partitioning assumes that the majority of states prior to the tipping point are reached by having fraction $c\in[0,1]$ bits flipped. The contribution of each system state over time, however, reflects a distribution of different states; reaching the tipping point from the ground state 0, can be done at $t-2$ prior to tipping by either remaining in 0.4 bits, or transitioning from 0.3 bits flipped to 0.4 and eventually to 0.5 in 2 time steps. The effect of these additional paths showed marginal effects on the integrated mutual information and asymptotic information.
Information flows conditioned on a partition is a form of conditional mutual information [James2016]. Prior results showed that conditional information produces synergy, i.e. information that is only present in the joint of all variables but cannot be found in any of the subset of each variable. Unfortunately, there is no generally agreed upon definition on how to measure synergy [Beer2015, Kolchinsky2022] and different estimates exist that may over or underestimate the synergetic effects. By partitioning one can create synergy as for a given partition each spin has some additional information about the other spins. For example, by taking the states such that $\langle S\rangle=0.1$, each spin “knows” that the average of the system equals 0.1. This creates shared information among the spins. Analyses were performed to estimate synergy using the redundancy estimation $I_{min}$[Williams2010]. Using this approach, no synergy was measured that affected the outcome of this study. However, it should be emphasized that synergetic effects may influence the causal interpretation of the approach presented here.
A general class of systems was studied governed by the Boltzmann-Gibbs distribution. For practical purposes the kinetic Ising model was only tested, but we speculate that the results should hold (in principle) for other systems dictated by the Boltzmann-Gibbs distribution. We leave the extension to other system Hamiltonians for future work.
7 Competing interests
The authors declare no competing interests.
8 Funding
This research is supported by grant Hyperion 2454972 of the Dutch National Police.
Appendix A Appendix
A.1 Background, scope & innovation
Noise induced transitions produces produces metastable behavior that is fundamental for the functioning of complex dynamical systems. For example, in neural systems, the presence of noise increases information processing. Similarly, the relation between glacial ice ages and earth eccentricity has been shown to have a strong correlation. Metastability manifests itself by means of noise that can be of two kinds [Forgoston2018]. External noise originates from events outside the internal system dynamics [Calim2021, Czaplicka2013a]. Examples include the influence of climate effects, population growth or a random noise source on a transmission line. External noise is commonly modeled by replacing an external control or order parameter by a stochastic process. Internal noise, in contrast, is inherent to the system itself and is caused by random interactions of elements of the system, e.g. individuals in a population, or molecules in chemical processes. Both types of noise can generate transitions between one metastable state and another. In this paper, the metastable behavior is studied of internal noise in complex dynamical networks governed by the kinetic Ising dynamics.
The ubiquity of multistability in complex systems calls for a general framework to understand how metastable transitions occur. The diversity of complex systems can be captured by an interaction networks that dynamically evolves over time. These dynamics can be seen as a distributive network of computational units, where each unit or element of the interaction network changes it state based on the input it gets from its local neighborhood. Lizier proposed that these proposed that the dynamic interaction of complex systems can be understood by their local information processing [Lizier2008, Lizier2013, Lizier2018]. Instead of describing the dynamics of the system in terms of their domain knowledge such as voltage over distance, disease spreading rate, or climate conditions, one can understand the dynamics in terms of the information dynamics. In particular, the field of information dynamics is concerned with describing the system behavior along its capacity to store information, transmit information, and modify information. By abstracting away the domain details of a system and recasting the dynamics in terms of how the system computes its next state, one can capture the intrinsic computation a system performs. The system behavior is encoded in terms of probability, and the relationship among these variables are explored using the language of information theory [Quax2017].
Information theory offers profound benefits over traditional methods used in meta-stability analysis as the methods developed are model-free, can capture non-linear relationships, can be used for both discrete and continuous variables, and can be estimated directly from data [Cover2005]. Shannon information measures such as mutual information and as well as Fisher information can be used to study how much information the system dynamics shares with the control parameter [Nicolis2016, Lizier2010].
Past research on information flows and metastable transitions focuses on methods to detect the onset of a tipping point [Scheffer2009, Prokopenko2011, Scheffer2001]. It often centers around an observation that the system’s ability to absorb noise reduces prior to the system going through a critical point. This critical slowing down, can be captured as a statistical signature where the Fisher information peaks [Eason2014]. However, these methods traditionally use some form of control parameter driving the system towards or away from a critical point. Most real-world systems lack such an explicit control parameter and require different methods. Furthermore, detecting a tipping point does not necessarily lead to further understanding how the tipping point was created. For example, for a finite size Ising model, the system produces bistable behavior. As one increases the noise parameter, the bistable behavior disappears. The increase in noise effectively changes the energy landscape, but little information is gained as to how initially the metastable behavior emerged.
In this work, a novel approach using information theory is explored to study metastable behavior. The statistical coherence between parts of the system are quantified by the the capability of individual nodes to predict the future behavior of the system [Lizier2013]. Two information features are introduced. Integrated mutual information measure predictive information of a node on the future of the system. Asymptotic information measures the long timescale memory capacity of a node. These measures differ from previous information methods such as transfer entropy [Schreiber], conditional mutual information under causal intervention [Ay2008], causation entropy [Runge2019], and time-delayed variants [Li2008] in that these methods are used to infer the transfer of information between sets of nodes by possible correcting for a third variable. Here, instead, we aim to understand how the elements in the system contribute to the macroscopic properties of the system. It is important to emphasize that information flows are not directly comparable to causal flows [James2016]. A rule of thumb is that causal flows focus on micro-level dynamics ($X$ causes $Y$), whereas information flows focus on the predictive aspects, a holistic view of emergent structures [Lizier2013]. In this sense, this work is similar to predictive information [Bialek1999] where predictive information of some system $S$ is projected onto its consistent elements $s_{i}\in S$ and computed as a function of time $t$.
A.2 Methods and definitions
A.2.1 Model
To study metastable behavior, we consider a system as a collection of random variables $S=\{s_{1},\dots,s_{n}\}$ governed by the Boltzmann-Gibbs distribution
$p(S)=\frac{1}{Z}\exp(-\beta\mathcal{H}(S)),$
where is the inverse temperature $\beta=\frac{1}{T}$ which control the noise in the system, $\mathcal{H}(S)$ is the system Hamiltonian which encodes the node-node dynamics. The choice of the energy function dictates what kind of system behavior we observe. Here, we focus on arguable the simplest models that shows metastable behavior: the kinetic Ising model, and the Susceptible-Infected-Susceptible model.
Temporal dynamics are simulated using Glauber dynamics sampling. In each discrete time step a spin is randomly chosen and a new state $X^{\prime}\in S$ is accepted with probability
$\begin{split}p(\text{accept}X^{\prime})=\frac{1}{1+\exp(-\beta\Delta E)},\end{split}$ (6)
where $\Delta E=\mathcal{H}(X^{\prime})-\mathcal{H}(X)$ is the energy difference between the current state $X$ and the proposed state $X^{\prime}$.
A.2.2 Kinetic Ising model
The traditional Ising model was originally developed to study ferromagnetism, and is considered one of the simplest models that generate complex behavior. It consists of a set of binary distributed spins $S=\{s_{1},\dots s_{n}\}$. Each spin contains energy given by the Hamiltonian
$\begin{split}\mathcal{H}(S)=-\sum_{i,j}J_{ij}s_{i}s_{j}-h_{i}s_{i}.\end{split}$ (7)
where $J_{ij}$ is the interaction energy of the spins $s_{i},s_{j}$.
The interaction energy effectively encodes the underlying network structure of the system. Different network structures are used in this study to provide a comprehensive numerical overview of the relation between network structure and information flows (see A.2). The interaction energy $J_{ij}$ is set to 1 if a connection exists in the network.
For sufficiently low noise (temperature), the Ising model shows metastable behavior (fig. 1c). Here, we aim to study how the system goes through a tipping point by tracking the information flow per node with the entire system state.
A.3 Information flow on complex networks
Informally, the information flows measures the statistical coherence between two random variables $X$ and $Y$ over time such that the present information in $Y$ cannot be explained by the past of $Y$ but rather by the past of $X$. Estimating information flow is inherently difficult due to the presence of confounding which potential traps the interpretation in the “correlation does not equal causation”. Under some context, however, information flow can be interpreted as causal [vanElteren2022]. Let $S=\{s_{1},\dots,s_{n}\}$ be a random process, and $S^{t}$ represent the state of the random process at some time $t$. The information present in $S$ is given as the Shannon entropy
$\begin{split}H(S)=-\sum_{x\in S}p(x)\log p(x)\end{split}$ (8)
where $\log$ is base 2 unless otherwise stated, and $p(x)$ is used as a short-hand for $p(S=x)$. Shannon entropy captures the uncertainty of a random variable; it can be understood as the number of yes/no questions needed to determine the state of $S$. This measure of uncertainty naturally extends to two variables with Shannon mutual information. Let $s_{i}$ be an element of the state of $S$, then the Shannon mutual information $I(S;s_{i})$ is given as
$\begin{split}I(S;s_{i})&=\sum_{S_{i}\in S,s^{\prime}\in s_{i}}p(S_{i},s^{\prime})\log\frac{p(S_{i},s^{\prime})}{p(S_{i})p(s^{\prime})}\\ &=H(S)-H(S|s_{i})\end{split}$ (9)
Shannon mutual information can be interpreted as the uncertainty reduction of $S$ after knowing the state of $s_{i}$. Consequently, it encodes how much statistical coherence $s_{i}$ and $S$ share. Shannon mutual information can be measured over time to encode how much information (in bits) flows from state $s_{i}^{\tau}$ to $S^{\tau+t}$
$\begin{split}I(S^{\tau+t};s_{i}^{\tau})=H(S^{\tau+t})-H(S^{\tau+t}|s_{i}^{\tau}).\end{split}$ (10)
Prior results showed that the nodes with the highest causal importance are those nodes that have the highest information flow (i.e. maximize 10) [vanElteren2022]. Intuitively, the nodes for which the future system “remembers” information from a node in the past, is the one that “drives” the system dynamics. Formally, these driver nodes can be identified by computing the total information flow between $S^{t}$ and $s_{i}$ can be captured with the integrated mutual information [vanElteren2022]
$\begin{split}\mu(s_{i})=\sum_{\tau=0}^{\infty}I(s_{i}^{t-\tau};S^{t}).\end{split}$ (11)
In some context, the nodes that maximizes the (11) are those nodes that have the highest causal influence in the system [vanElteren2022]. However in general information flows are difficult to equate to causal flows [Lizier2013, James2016]. Here, the local information flows are computed by considering the integrated mutual information conditioned on part of the entire state space. This allows for mapping the local information flows between nodes and the system over time, but does not guarantee that the measured information flows are directly causal. The main reason being that having predictive power about the future, could be completely caused by the partitioning. In [vanElteren2022] the correlation measured considered all possible states, and the measures were directly related to a causal effect.
In addition, in [vanElteren2022] the shared information between the system with a node shifted over time ($I(S^{\tau}:s_{i}^{\tau+t})$) was considered. Applying this approach under a state partition $I(S^{\tau}:s_{i}^{\tau+t}|\langle S\rangle)$causes a violation of the data processing as information may flow from a node at a particular $t=t_{1}$ and then flow back to the node at $t=t2,t_{2}>t_{1}$. In order to simplify the interpretation of the information flows and keep the data processing inequality, the reverse $I(S^{t+\tau}:s_{i}^{\tau}|\langle S\rangle)$ was computed in the present study.
A.4 Noise matching procedure
The Boltzmann-Gibbs distribution is parameterized by noise factor $\beta=\frac{1}{kT}$ where $T$ is the temperature and $k$ is the Boltzmann constant. For high $\beta$ values metastable behavior occurs in the kinetic Ising model. The temperature was chosen such that the statistical complexity [Lopez-Ruiz1995a] was maximized. The statistical complexity $C$ is computed as
$C=\bar{H}(S)D(S),$
where $\bar{H}(S)=\frac{H(s)}{-\log_{2}(|S|)}$ is the system entropy, and $D(S)$ measures the distance to disequilibrium
$D(S)=\sum_{i}(p(S_{i})-\frac{1}{|S|})^{2}.$
A typical statistical complexity curve is seen in fig. A.6. The noise parameter $\beta$ is set such that it maximizes the statistical complexity using numerical optimization (COBYLA method in scipy’s optimize.minimize module) [Virtanen2020].
Figure A.6: (a) Statistical complexity ($C$), normalized system entropy ($H(S)$) and disequilibrium ($D(S)$) as a function of the temperature ($T=\frac{1}{\beta}$) for Krackhardt kite graph. The noise parameter was set such that it maximizes the statistical complexity (vertical black line). The values are normalized between [0,1] for aesthetic purposes. (b) State distribution $p(S)$ for temperature that maximizes the statistical complexity in (a) as a function of nodes in state 1.
A.5 Exact information flows $I(s_{i}^{\tau};S^{\tau+t})$
In order to compute $I(s_{i}^{\tau}:S^{\tau+t})$, the conditional distribution $p(S^{\tau+t}|s_{i}^{\tau})$ and $p(S^{\tau+t})$ needs to be computed. For Glauber dynamics, the system $S$ transitions into $S^{\prime}$ by considering to flips by randomly choosing node $s_{i}$. The transition matrix $p(S^{t}|s_{i})=\textbf{P}$ can be constructed by computing each entry $p_{ij}$ as
$\begin{split}p_{ij,i\neq j}&=\frac{1}{|S|}\frac{1}{1+\exp(-\Delta E)}\\ p_{ii}&=1-\sum_{j,j\neq i}p_{ij},\end{split}$
where $\Delta E=\mathcal{H}(S_{j})-\mathcal{H}(S_{j})$ encodes the energy difference of moving from $S_{i}$ to $S_{j}$. The state to state transition P matrix will be of size $2^{|S|}\times 2^{|S|}\times|\mathcal{A}_{s_{i}}|$, where $|\mathcal{A}_{s_{i}}|$ is the size of the alphabet of $s_{i}$, which becomes computationally intractable due to its exponential growth with the system size $|S|$. The exact information flows can then be computed by evaluating $p(S^{t}|s_{i})$ out of equilibrium by evaluating all $S^{t}$ for all possible node states $s_{i}$ where $p(S^{t})$ is computed as
$p(S^{\tau+t})=\sum_{s_{i}}p(S^{\tau+t}|s_{i}^{\tau})p(s_{i}^{\tau}).$
A.5.1 Extrapolation with regressions
Exact information flows were computed per graph for $t=500$ times steps. Using ordinary least squares a double exponential was fit to estimate the information flows for longer $t$ and estimate the integrated mutual information and asymptotic information.
A.6 Noise estimation procedure
Tipping point behavior under intervention was quantified by evaluating the level of noise on both side of the tipping point. Let $T1$ represent the ground state where all spins are 0, $T2$ where all spins, and the tipping point $TP$ is where the instantaneous macrostate $M(S^{t})=0.5$. Fluctuations of the system macrostate was evaluated by analyzing the second moment above and below the tipping point. This was achieved by numerically simulating the system trajectories under 6 different seeds for $t=10e6$ time-steps. The data was split between two sets (above and below the tipping point) and the noise $\eta$ was computed as
$\begin{split}\eta=\frac{1}{\alpha^{2}|S_{w}|}\sum_{w}{S_{w}^{t}}^{2},\end{split}$
where $w\in\{\langle S\rangle<0.5,\langle S\rangle>0.5\}$, and
$S_{w}^{t}=\Bigl{\{}\begin{aligned} S^{t}&\textrm{ if }S^{t}<0.5\\ 1-S^{t}&\textrm{ if }S^{t}>0.5\end{aligned}$ (12)
is the instantaneous system trajectory for the system macrostate above or below the tipping point value. The factor $\alpha$ corrects for the reduced range the system macrostate has under interventions. For example pinning a node $s_{i}$ to state 0, reduces the maximum possible macrostate to $1-\frac{1}{n}$ where $n$ is the size of the system. The correction factor $\alpha$ is set such that for an intervention on 0 for a particular node, the range $S_{\langle S\rangle>0.5}$ alpha is set to $\frac{n}{2}-\frac{1}{n}$.
A.7 Switch susceptibility as a function of degree
First, we investigate the susceptibility of a spin as a function of its degree. The susceptibility of a spin switching its state is a function both of the system temperature $T$ and the system dynamics. The system dynamics would contribute to the susceptibility through the underlying network structure either directly or indirectly. The network structure produces local correlations which affects the switch probability for a given spin.
As an initial approximation, we consider the susceptibility of a target spin $s_{i}$ to flip from a majority state to a minority state given the state of its neighbors where the neighbors are not connected among themselves. Further, the assumption is that for the instantaneous update of $s_{i}$ the configuration of the neighborhood of $s_{i}$ can be considered as the outcome of a binomial trial. Let, $N$ be a random variable with state space $\{0,1\}^{|N|}$, and let $n_{j}\in N$ represent a neighbor of $s_{i}$. We assume that all neighbors of $s_{i}$ are i.i.d. distributed given the instantaneous system magnetization
$M(S^{t})=\frac{1}{|S^{t}|}\sum_{i}s_{i}^{t}.$
Let the minority state be 1 and the majority state be 0, the expectation of $s_{i}$ flipping from the majority state to the minority state is given as:
$E[p(s_{i}=1|N)]_{p(N)}=\sum_{N_{i}\in N}p(N_{i})p(s_{i}=1|N_{i})\\ =\sum_{N_{i}\in N}\prod_{j}^{|N_{i}|}p(n_{j})p(s_{i}=1|N_{i})\\ =\sum_{N_{i}\in N}{n\choose k}f^{k}(1-f)^{n-k}p(s_{i}=1|f),$ (13)
where $f$ is the fraction of nodes in the majority states, $n$ is the number of neighbors, $k$ is the number of nodes in state 0. In fig. A.8. This is computed as a function of the degree of spin $s_{i}$. As the degree increases, the susceptibility for a spin decreases relatively to the same spin with a lower degree. This implies that the susceptibility of change to random fluctuations are more likely to occur in nodes with less external constraints as measured by degree.
A.8 Additional networks
The kite graph was chosen as it allowed for computing exact information flows while retaining a high variety of degree distribution given the small size. Other networks were also tested. In fig. A.7) different network structure were used. Each node is governed by kinetic Ising spin dynamics.
Figure A.7: Adjusted mutual information for a random tree (top), and Leder-Coxeter Fruchte graphs (middle, bottom). Each node is goverened by kinetic Ising spin dyanmics. Far away from the tipping point (fraction nodes 1 = 0.5) most information flows are concentrated on non-hub nodes. As the system approaches the tipping point (fraction = 0.5), the information flows move inwards, generating higher adjusted integrated mutual information for nodes with higher degree.
A.9 Flip probability per degree
In fig. A.8 the tendency for a node to flip from the majority to the minority state is computed as function of fraction of nodes possessing the majority states 1 in the system, denoted as $N$. Two things are observed. First, nodes with lower degree are more susceptible to noise than nodes with higher degree. For a given system stability, nodes with lower degree tend to have a higher tendency to flip. This is true for all distances of the system to the tipping point. In contrast, the higher the degree of the node, the closer the system has to be to a tipping point for the node to change its state. This can be explained by the fact that lower degree nodes, have fewer constraints compared to nodes with higher degree nodes. For Ising spin kinetics, the nodes with higher degree tend to be more “frozen” in their node dynamics than nodes with lower degree. Second, in order for a node to flip with probability with similar mass, i.e. ($E[p(s_{i})|N]=0.2$) a node with higher degree needs to be closer to the tipping point than nodes with lower degree. In fact, the order of susceptibility is correlated with the degree; the susceptibility decreases with increasing degree and fixed fraction of nodes in state 1.
Figure A.8: Susceptibility of a node with degree $k$ switching from the minority state 0 to the majority state 1 as a function of the neighborhood entropy for $\beta=0.5$. The neighborhood entropy encodes how stable the environment of a spin is. As the system approaches the tipping point, the propensity of a node to flip from to the minority state increases faster for low degree nodes than for high degree nodes. Higher degree nodes require more change in their local environment to flip to the majority state. See for details A.7.
Figure A.9: Shortest path analysis of the system ending up in the tipping point from the state where all nodes have state 0. The node size is proportional to the expectation value of a node having state 1 ($E[s_{i}=1]_{S^{t},M(S^{5})}$ as a function of the fraction of nodes having state 1. The expectation values are computed based on 30240 trajectories, an example trajectory can be seen in fig. 3.
A.10 Synthetic networks
For the synthetic graphs, 100 non-isomorphic connected Erdos-Renyi networks were generated with a p = 0.2. Graphs were generated randomly and rejected if the graph did not contain a giant component, or was isomorphic with already generated graphs. For each of the graphs, information curves were computed as function of the macrostate $\langle S\rangle$.
Figure A.10: Erdos-Renyi graphs generated from seed = 0 to produce non-isomorphic connected graphs.
A.10.1 Noise and time spent
Various network structures are generated in the synthetic networks. The variety of network structure has non-linear effects on the information flows. The effect of intervention in fig. 5 is made relative to the control values for the graph and seed. The second moment (appendix: A.6) and the time spent below the tipping point are normalized with respect to the graph (fig. A.10) and the seed. In total 6 seeds are used (0, 12, 123, 1234, 123456, 1234567).

### v3

@@
 1 Introduction
-Multistability is a crucial characteristic observed in various real-world complex systems [Ladyman2013, vanNes2016].
+Multistability is a crucial characteristic observed in various real-world complex systems [1, 2].
 It describes the phenomenon in which a system, under the influence of noise, explores different states within its state space at various spatio-temporal scales.
-Examples range from the brain’s ability to operate in one cognitive mode for an extended period before rapidly changing its state [Kandel2000, Fries2015].
-Larger-scale instances include multistability in opinion dynamics [Galam2020] and ecosystems or climate systems balancing between metastable states [1986, Wunderling2021].
-Evidence increasingly suggests that noise fluctuations play a fundamental role in transitions between attractor states [Beggs2012, Mitchell1993, Mitchella, Forgoston2018a].
+Examples range from the brain’s ability to operate in one cognitive mode for an extended period before rapidly changing its state [3, 4].
+Larger-scale instances include multistability in opinion dynamics [5] and ecosystems or climate systems balancing between metastable states [6, 7].
+Evidence increasingly suggests that noise fluctuations play a fundamental role in transitions between attractor states [8, 9, 10, 11].
 Traditionally, noise is perceived as an unwanted signal, diminishing system effectiveness or signal quality.
-However, in nonlinear systems, such as interconnected neurons in the brain, noise can play a surprising role by enabling the exploration of larger parts of the state space, allowing the system to escape local minima [Czaplicka2013, Nicolis2016].
+However, in nonlinear systems, such as interconnected neurons in the brain, noise can play a surprising role by enabling the exploration of larger parts of the state space, allowing the system to escape local minima [12, 13].
 Here, we focus on a particular subset of dynamical systems where each node’s state in the next time step is governed by the Boltzmann-Gibbs distribution, akin to a local thermal equilibrium perspective.
-Such models have been applied to describe various behaviors, including neural dynamics [Hopfield1982b], opinion dynamics, ferromagnetic spins [Glauber1963], and organized criminal gang interactions [DOrsogna2015a].
+Such models have been applied to describe various behaviors, including neural dynamics [14], opinion dynamics, ferromagnetic spins [15], and organized criminal gang interactions [16].
 Dynamical networks governed by the Boltzmann-Gibbs distribution involve nodes choosing a new state based on the local equilibrium potential induced by their neighbors.
@@
 A global ‘temperature’ parameter introduces noise, influencing the likelihood of choosing higher energy states.
-For networked dynamical system, the structure itself significantly influences systemic transitions, making them more likely than in systems without network effects [Harush2017a, Gao2016, Wunderling2020, Wunderling2021].
+For networked dynamical system, the structure itself significantly influences systemic transitions, making them more likely than in systems without network effects [17, 18, 19, 7].
 Our method uncovers the network percolation process facilitating endogenously generated, noise-induced transitions, offering a different perspective from traditional approaches.
@@
 These measures, namely integrated mutual information and asymptotic information, are designed to capture the effects of short-term and long-term correlations, respectively.
-They provide valuable insights into the dynamics of metastable transitions within networked systems and can be directly computed on observations.
+They provide valuable insights into the dynamics of metastable transitions within networked systems and can be directly computed from observations.
 Figure 1: A dynamical network governed by kinetic Ising dynamics produces multistable behavior.
@@
 2 Methods
-Fluctuations and their correlations at time $\tau$ are captured using Shannon’s mutual information [Cover2005] shared between a node’s state ($s_{i}^{t}$) at time $t$ and the entire future system state ($S^{t+\tau}$).
-$I(s_{i}^{\tau}:S^{\tau+t})$.
+Fluctuations and their correlations at time $\tau$ are captured using Shannon’s mutual information [20] shared between a node’s state ($s_{i}^{t}$) at time $t$ and the entire future system state ($S^{t+\tau}$) $I(s_{i}^{\tau}:S^{\tau+t})$.
 The time lag $t$ is used to analyze two key features of information flows of a system: the area under the curve (AUC) of short-term information, and sustained level of long term information.
 The contribution of a node to the dynamics of the system will differ depending on the network connectivity of a node (fig.
-A.8) [vanElteren2022, Quax2013].
+A.8) [21, 22].
 The total amount of fluctuations shared between the node’s current state and the system’s short-term future trajectory is computed as the integrated mutual information
 $\begin{split}\mu(s_{i})=\sum_{t=0}^{\infty}(I(s_{i}^{\tau}:S^{\tau+t})-\omega_{s_{i}})\Delta t.\end{split}$ (2)
-Intuitively, $\mu(s_{i})$ represents a combination of the intensity and duration of the short-term fluctuations on the (transient) system dynamics [vanElteren2022].
+Intuitively, $\mu(s_{i})$ represents a combination of the intensity and duration of the short-term fluctuations on the (transient) system dynamics [21].
 It reflects how much of the node state is in the “working memory” of the system.
@@
 However, these nodes lack long-term predictive information about future system states.
-Conversely, a node classified as -1, has is a pure stabilizer.
+Conversely, a node classified as -1, is a pure stabilizer.
 Roles having a value $r_{i}\sim 0$ are more difficult to interpret as the zero value could be caused by an equally large integrated mutual information and asymptotic information or a generally lacking high score in both.
@@
 Due to their high degree of freedom, these nodes are more effected by external noise.
-They are instigators and inject propagate in the system, destabilizing more stable nodes.
+They are instigators and propagate noise in the system, destabilizing more stable nodes.
 In contrast, stabilizer nodes, have low degree of freedom and require more energy to change state.
@@
 On rare occasions, however, the cascade is propagated progressively from low degree, to higher and higher degree.
-A similar domino mechanism was recently found in climate science [Wunderling2020, Wunderling2021].
+A similar domino mechanism was recently found in climate science [19, 7].
 Wunderling and colleagues provided a simplified model of the climate system, analyzing how various components contribute to the stability of the climate.
@@
 The effect of these additional paths showed marginal effects on the integrated mutual information and asymptotic information.
-Information flows conditioned on a partition is a form of conditional mutual information [James2016].
+Information flows conditioned on a partition is a form of conditional mutual information [23].
 Prior results showed that conditional information produces synergy, i.e.
 information that is only present in the joint of all variables but cannot be found in any of the subset of each variable.
-Unfortunately, there is no generally agreed upon definition on how to measure synergy [Beer2015, Kolchinsky2022] and different estimates exist that may over or underestimate the synergetic effects.
+Unfortunately, there is no generally agreed upon definition on how to measure synergy [24, 25] and different estimates exist that may over or underestimate the synergetic effects.
 By partitioning one can create synergy as for a given partition each spin has some additional information about the other spins.
@@
 This creates shared information among the spins.
-Analyses were performed to estimate synergy using the redundancy estimation $I_{min}$[Williams2010].
+Analyses were performed to estimate synergy using the redundancy estimation $I_{min}$[26].
 Using this approach, no synergy was measured that affected the outcome of this study.
@@
 Similarly, the relation between glacial ice ages and earth eccentricity has been shown to have a strong correlation.
-Metastability manifests itself by means of noise that can be of two kinds [Forgoston2018].
-External noise originates from events outside the internal system dynamics [Calim2021, Czaplicka2013a].
+Metastability manifests itself by means of noise that can be of two kinds [27].
+External noise originates from events outside the internal system dynamics [28, 29].
 Examples include the influence of climate effects, population growth or a random noise source on a transmission line.
@@
 These dynamics can be seen as a distributive network of computational units, where each unit or element of the interaction network changes it state based on the input it gets from its local neighborhood.
-Lizier proposed that these proposed that the dynamic interaction of complex systems can be understood by their local information processing [Lizier2008, Lizier2013, Lizier2018].
+Lizier proposed that these proposed that the dynamic interaction of complex systems can be understood by their local information processing [30, 31, 32].
 Instead of describing the dynamics of the system in terms of their domain knowledge such as voltage over distance, disease spreading rate, or climate conditions, one can understand the dynamics in terms of the information dynamics.
@@
 By abstracting away the domain details of a system and recasting the dynamics in terms of how the system computes its next state, one can capture the intrinsic computation a system performs.
-The system behavior is encoded in terms of probability, and the relationship among these variables are explored using the language of information theory [Quax2017].
-Information theory offers profound benefits over traditional methods used in meta-stability analysis as the methods developed are model-free, can capture non-linear relationships, can be used for both discrete and continuous variables, and can be estimated directly from data [Cover2005].
-Shannon information measures such as mutual information and as well as Fisher information can be used to study how much information the system dynamics shares with the control parameter [Nicolis2016, Lizier2010].
-Past research on information flows and metastable transitions focuses on methods to detect the onset of a tipping point [Scheffer2009, Prokopenko2011, Scheffer2001].
+The system behavior is encoded in terms of probability, and the relationship among these variables are explored using the language of information theory [33].
+Information theory offers profound benefits over traditional methods used in meta-stability analysis as the methods developed are model-free, can capture non-linear relationships, can be used for both discrete and continuous variables, and can be estimated directly from data [20].
+Shannon information measures such as mutual information and as well as Fisher information can be used to study how much information the system dynamics shares with the control parameter [13, 34].
+Past research on information flows and metastable transitions focuses on methods to detect the onset of a tipping point [35, 36, 37].
 It often centers around an observation that the system’s ability to absorb noise reduces prior to the system going through a critical point.
-This critical slowing down, can be captured as a statistical signature where the Fisher information peaks [Eason2014].
+This critical slowing down, can be captured as a statistical signature where the Fisher information peaks [38].
 However, these methods traditionally use some form of control parameter driving the system towards or away from a critical point.
@@
 In this work, a novel approach using information theory is explored to study metastable behavior.
-The statistical coherence between parts of the system are quantified by the the capability of individual nodes to predict the future behavior of the system [Lizier2013].
+The statistical coherence between parts of the system are quantified by the the capability of individual nodes to predict the future behavior of the system [31].
 Two information features are introduced.
@@
 Asymptotic information measures the long timescale memory capacity of a node.
-These measures differ from previous information methods such as transfer entropy [Schreiber], conditional mutual information under causal intervention [Ay2008], causation entropy [Runge2019], and time-delayed variants [Li2008] in that these methods are used to infer the transfer of information between sets of nodes by possible correcting for a third variable.
+These measures differ from previous information methods such as transfer entropy [39], conditional mutual information under causal intervention [40], causation entropy [41], and time-delayed variants [42] in that these methods are used to infer the transfer of information between sets of nodes by possible correcting for a third variable.
 Here, instead, we aim to understand how the elements in the system contribute to the macroscopic properties of the system.
-It is important to emphasize that information flows are not directly comparable to causal flows [James2016].
-A rule of thumb is that causal flows focus on micro-level dynamics ($X$ causes $Y$), whereas information flows focus on the predictive aspects, a holistic view of emergent structures [Lizier2013].
-In this sense, this work is similar to predictive information [Bialek1999] where predictive information of some system $S$ is projected onto its consistent elements $s_{i}\in S$ and computed as a function of time $t$.
+It is important to emphasize that information flows are not directly comparable to causal flows [23].
+A rule of thumb is that causal flows focus on micro-level dynamics ($X$ causes $Y$), whereas information flows focus on the predictive aspects, a holistic view of emergent structures [31].
+In this sense, this work is similar to predictive information [43] where predictive information of some system $S$ is projected onto its consistent elements $s_{i}\in S$ and computed as a function of time $t$.
 A.2 Methods and definitions
@@
 Estimating information flow is inherently difficult due to the presence of confounding which potential traps the interpretation in the “correlation does not equal causation”.
-Under some context, however, information flow can be interpreted as causal [vanElteren2022].
+Under some context, however, information flow can be interpreted as causal [21].
 Let $S=\{s_{1},\dots,s_{n}\}$ be a random process, and $S^{t}$ represent the state of the random process at some time $t$.
@@
 Prior results showed that the nodes with the highest causal importance are those nodes that have the highest information flow (i.e.
-maximize 10) [vanElteren2022].
+maximize 10) [21].
 Intuitively, the nodes for which the future system “remembers” information from a node in the past, is the one that “drives” the system dynamics.
-Formally, these driver nodes can be identified by computing the total information flow between $S^{t}$ and $s_{i}$ can be captured with the integrated mutual information [vanElteren2022]
+Formally, these driver nodes can be identified by computing the total information flow between $S^{t}$ and $s_{i}$ can be captured with the integrated mutual information [21]
 $\begin{split}\mu(s_{i})=\sum_{\tau=0}^{\infty}I(s_{i}^{t-\tau};S^{t}).\end{split}$ (11)
-In some context, the nodes that maximizes the (11) are those nodes that have the highest causal influence in the system [vanElteren2022].
-However in general information flows are difficult to equate to causal flows [Lizier2013, James2016].
+In some context, the nodes that maximizes the (11) are those nodes that have the highest causal influence in the system [21].
+However in general information flows are difficult to equate to causal flows [31, 23].
 Here, the local information flows are computed by considering the integrated mutual information conditioned on part of the entire state space.
@@
 The main reason being that having predictive power about the future, could be completely caused by the partitioning.
-In [vanElteren2022] the correlation measured considered all possible states, and the measures were directly related to a causal effect.
-In addition, in [vanElteren2022] the shared information between the system with a node shifted over time ($I(S^{\tau}:s_{i}^{\tau+t})$) was considered.
+In [21] the correlation measured considered all possible states, and the measures were directly related to a causal effect.
+In addition, in [21] the shared information between the system with a node shifted over time ($I(S^{\tau}:s_{i}^{\tau+t})$) was considered.
 Applying this approach under a state partition $I(S^{\tau}:s_{i}^{\tau+t}|\langle S\rangle)$causes a violation of the data processing as information may flow from a node at a particular $t=t_{1}$ and then flow back to the node at $t=t2,t_{2}>t_{1}$.
@@
 For high $\beta$ values metastable behavior occurs in the kinetic Ising model.
-The temperature was chosen such that the statistical complexity [Lopez-Ruiz1995a] was maximized.
+The temperature was chosen such that the statistical complexity [44] was maximized.
 The statistical complexity $C$ is computed as
@@
 A.6.
-The noise parameter $\beta$ is set such that it maximizes the statistical complexity using numerical optimization (COBYLA method in scipy’s optimize.minimize module) [Virtanen2020].
+The noise parameter $\beta$ is set such that it maximizes the statistical complexity using numerical optimization (COBYLA method in scipy’s optimize.minimize module) [45].
 Figure A.6: (a) Statistical complexity ($C$), normalized system entropy ($H(S)$) and disequilibrium ($D(S)$) as a function of the temperature ($T=\frac{1}{\beta}$) for Krackhardt kite graph.

### v4

@@
-Cascades towards noise-induced transitions on networks revealed using information flows
+./main.bib
 Abstract
@@
 1 Introduction
-Multistability is a crucial characteristic observed in various real-world complex systems [1, 2].
+Multistability is a crucial characteristic observed in various real-world complex systems [Ladyman2013, vanNes2016].
 It describes the phenomenon in which a system, under the influence of noise, explores different states within its state space at various spatio-temporal scales.
-Examples range from the brain’s ability to operate in one cognitive mode for an extended period before rapidly changing its state [3, 4].
-Larger-scale instances include multistability in opinion dynamics [5] and ecosystems or climate systems balancing between metastable states [6, 7].
-Evidence increasingly suggests that noise fluctuations play a fundamental role in transitions between attractor states [8, 9, 10, 11].
+Examples range from the brain’s ability to operate in one cognitive mode for an extended period before rapidly changing its state [Kandel2000, Fries2015].
+Larger-scale instances include multistability in opinion dynamics [Galam2020] and ecosystems or climate systems balancing between metastable states [1986, Wunderling2021].
+Evidence increasingly suggests that noise fluctuations play a fundamental role in transitions between attractor states [Beggs2012, Mitchell1993, Mitchella, Forgoston2018a].
 Traditionally, noise is perceived as an unwanted signal, diminishing system effectiveness or signal quality.
-However, in nonlinear systems, such as interconnected neurons in the brain, noise can play a surprising role by enabling the exploration of larger parts of the state space, allowing the system to escape local minima [12, 13].
+However, in nonlinear systems, such as interconnected neurons in the brain, noise can play a surprising role by enabling the exploration of larger parts of the state space, allowing the system to escape local minima [Czaplicka2013, Nicolis2016].
 Here, we focus on a particular subset of dynamical systems where each node’s state in the next time step is governed by the Boltzmann-Gibbs distribution, akin to a local thermal equilibrium perspective.
-Such models have been applied to describe various behaviors, including neural dynamics [14], opinion dynamics, ferromagnetic spins [15], and organized criminal gang interactions [16].
+Such models have been applied to describe various behaviors, including neural dynamics [Hopfield1982b], opinion dynamics, ferromagnetic spins [Glauber1963], and organized criminal gang interactions [DOrsogna2015a].
 Dynamical networks governed by the Boltzmann-Gibbs distribution involve nodes choosing a new state based on the local equilibrium potential induced by their neighbors.
@@
 A global ‘temperature’ parameter introduces noise, influencing the likelihood of choosing higher energy states.
-For networked dynamical system, the structure itself significantly influences systemic transitions, making them more likely than in systems without network effects [17, 18, 19, 7].
+For networked dynamical system, the structure itself significantly influences systemic transitions, making them more likely than in systems without network effects [Harush2017a, Gao2016, Wunderling2020, Wunderling2021].
 Our method uncovers the network percolation process facilitating endogenously generated, noise-induced transitions, offering a different perspective from traditional approaches.
@@
 2 Methods
-Fluctuations and their correlations at time $\tau$ are captured using Shannon’s mutual information [20] shared between a node’s state ($s_{i}^{t}$) at time $t$ and the entire future system state ($S^{t+\tau}$) $I(s_{i}^{\tau}:S^{\tau+t})$.
+Fluctuations and their correlations at time $\tau$ are captured using Shannon’s mutual information [Cover2005] shared between a node’s state ($s_{i}^{t}$) at time $t$ and the entire future system state ($S^{t+\tau}$) $I(s_{i}^{\tau}:S^{\tau+t})$.
 The time lag $t$ is used to analyze two key features of information flows of a system: the area under the curve (AUC) of short-term information, and sustained level of long term information.
 The contribution of a node to the dynamics of the system will differ depending on the network connectivity of a node (fig.
-A.8) [21, 22].
+A.8) [vanElteren2022, Quax2013].
 The total amount of fluctuations shared between the node’s current state and the system’s short-term future trajectory is computed as the integrated mutual information
 $\begin{split}\mu(s_{i})=\sum_{t=0}^{\infty}(I(s_{i}^{\tau}:S^{\tau+t})-\omega_{s_{i}})\Delta t.\end{split}$ (2)
-Intuitively, $\mu(s_{i})$ represents a combination of the intensity and duration of the short-term fluctuations on the (transient) system dynamics [21].
+Intuitively, $\mu(s_{i})$ represents a combination of the intensity and duration of the short-term fluctuations on the (transient) system dynamics [vanElteren2022].
 It reflects how much of the node state is in the “working memory” of the system.
@@
 On rare occasions, however, the cascade is propagated progressively from low degree, to higher and higher degree.
-A similar domino mechanism was recently found in climate science [19, 7].
+A similar domino mechanism was recently found in climate science [Wunderling2020, Wunderling2021].
 Wunderling and colleagues provided a simplified model of the climate system, analyzing how various components contribute to the stability of the climate.
@@
 The effect of these additional paths showed marginal effects on the integrated mutual information and asymptotic information.
-Information flows conditioned on a partition is a form of conditional mutual information [23].
+Information flows conditioned on a partition is a form of conditional mutual information [James2016].
 Prior results showed that conditional information produces synergy, i.e.
 information that is only present in the joint of all variables but cannot be found in any of the subset of each variable.
-Unfortunately, there is no generally agreed upon definition on how to measure synergy [24, 25] and different estimates exist that may over or underestimate the synergetic effects.
+Unfortunately, there is no generally agreed upon definition on how to measure synergy [Beer2015, Kolchinsky2022] and different estimates exist that may over or underestimate the synergetic effects.
 By partitioning one can create synergy as for a given partition each spin has some additional information about the other spins.
@@
 This creates shared information among the spins.
-Analyses were performed to estimate synergy using the redundancy estimation $I_{min}$[26].
+Analyses were performed to estimate synergy using the redundancy estimation $I_{min}$[Williams2010].
 Using this approach, no synergy was measured that affected the outcome of this study.
@@
 Similarly, the relation between glacial ice ages and earth eccentricity has been shown to have a strong correlation.
-Metastability manifests itself by means of noise that can be of two kinds [27].
-External noise originates from events outside the internal system dynamics [28, 29].
+Metastability manifests itself by means of noise that can be of two kinds [Forgoston2018].
+External noise originates from events outside the internal system dynamics [Calim2021, Czaplicka2013a].
 Examples include the influence of climate effects, population growth or a random noise source on a transmission line.
@@
 These dynamics can be seen as a distributive network of computational units, where each unit or element of the interaction network changes it state based on the input it gets from its local neighborhood.
-Lizier proposed that these proposed that the dynamic interaction of complex systems can be understood by their local information processing [30, 31, 32].
+Lizier proposed that these proposed that the dynamic interaction of complex systems can be understood by their local information processing [Lizier2008, Lizier2013, Lizier2018].
 Instead of describing the dynamics of the system in terms of their domain knowledge such as voltage over distance, disease spreading rate, or climate conditions, one can understand the dynamics in terms of the information dynamics.
@@
 By abstracting away the domain details of a system and recasting the dynamics in terms of how the system computes its next state, one can capture the intrinsic computation a system performs.
-The system behavior is encoded in terms of probability, and the relationship among these variables are explored using the language of information theory [33].
-Information theory offers profound benefits over traditional methods used in meta-stability analysis as the methods developed are model-free, can capture non-linear relationships, can be used for both discrete and continuous variables, and can be estimated directly from data [20].
-Shannon information measures such as mutual information and as well as Fisher information can be used to study how much information the system dynamics shares with the control parameter [13, 34].
-Past research on information flows and metastable transitions focuses on methods to detect the onset of a tipping point [35, 36, 37].
+The system behavior is encoded in terms of probability, and the relationship among these variables are explored using the language of information theory [Quax2017].
+Information theory offers profound benefits over traditional methods used in meta-stability analysis as the methods developed are model-free, can capture non-linear relationships, can be used for both discrete and continuous variables, and can be estimated directly from data [Cover2005].
+Shannon information measures such as mutual information and as well as Fisher information can be used to study how much information the system dynamics shares with the control parameter [Nicolis2016, Lizier2010].
+Past research on information flows and metastable transitions focuses on methods to detect the onset of a tipping point [Scheffer2009, Prokopenko2011, Scheffer2001].
 It often centers around an observation that the system’s ability to absorb noise reduces prior to the system going through a critical point.
-This critical slowing down, can be captured as a statistical signature where the Fisher information peaks [38].
+This critical slowing down, can be captured as a statistical signature where the Fisher information peaks [Eason2014].
 However, these methods traditionally use some form of control parameter driving the system towards or away from a critical point.
@@
 In this work, a novel approach using information theory is explored to study metastable behavior.
-The statistical coherence between parts of the system are quantified by the the capability of individual nodes to predict the future behavior of the system [31].
+The statistical coherence between parts of the system are quantified by the the capability of individual nodes to predict the future behavior of the system [Lizier2013].
 Two information features are introduced.
@@
 Asymptotic information measures the long timescale memory capacity of a node.
-These measures differ from previous information methods such as transfer entropy [39], conditional mutual information under causal intervention [40], causation entropy [41], and time-delayed variants [42] in that these methods are used to infer the transfer of information between sets of nodes by possible correcting for a third variable.
+These measures differ from previous information methods such as transfer entropy [Schreiber], conditional mutual information under causal intervention [Ay2008], causation entropy [Runge2019], and time-delayed variants [Li2008] in that these methods are used to infer the transfer of information between sets of nodes by possible correcting for a third variable.
 Here, instead, we aim to understand how the elements in the system contribute to the macroscopic properties of the system.
-It is important to emphasize that information flows are not directly comparable to causal flows [23].
-A rule of thumb is that causal flows focus on micro-level dynamics ($X$ causes $Y$), whereas information flows focus on the predictive aspects, a holistic view of emergent structures [31].
-In this sense, this work is similar to predictive information [43] where predictive information of some system $S$ is projected onto its consistent elements $s_{i}\in S$ and computed as a function of time $t$.
+It is important to emphasize that information flows are not directly comparable to causal flows [James2016].
+A rule of thumb is that causal flows focus on micro-level dynamics ($X$ causes $Y$), whereas information flows focus on the predictive aspects, a holistic view of emergent structures [Lizier2013].
+In this sense, this work is similar to predictive information [Bialek1999] where predictive information of some system $S$ is projected onto its consistent elements $s_{i}\in S$ and computed as a function of time $t$.
 A.2 Methods and definitions
@@
 Estimating information flow is inherently difficult due to the presence of confounding which potential traps the interpretation in the “correlation does not equal causation”.
-Under some context, however, information flow can be interpreted as causal [21].
+Under some context, however, information flow can be interpreted as causal [vanElteren2022].
 Let $S=\{s_{1},\dots,s_{n}\}$ be a random process, and $S^{t}$ represent the state of the random process at some time $t$.
@@
 Prior results showed that the nodes with the highest causal importance are those nodes that have the highest information flow (i.e.
-maximize 10) [21].
+maximize 10) [vanElteren2022].
 Intuitively, the nodes for which the future system “remembers” information from a node in the past, is the one that “drives” the system dynamics.
-Formally, these driver nodes can be identified by computing the total information flow between $S^{t}$ and $s_{i}$ can be captured with the integrated mutual information [21]
+Formally, these driver nodes can be identified by computing the total information flow between $S^{t}$ and $s_{i}$ can be captured with the integrated mutual information [vanElteren2022]
 $\begin{split}\mu(s_{i})=\sum_{\tau=0}^{\infty}I(s_{i}^{t-\tau};S^{t}).\end{split}$ (11)
-In some context, the nodes that maximizes the (11) are those nodes that have the highest causal influence in the system [21].
-However in general information flows are difficult to equate to causal flows [31, 23].
+In some context, the nodes that maximizes the (11) are those nodes that have the highest causal influence in the system [vanElteren2022].
+However in general information flows are difficult to equate to causal flows [Lizier2013, James2016].
 Here, the local information flows are computed by considering the integrated mutual information conditioned on part of the entire state space.
@@
 The main reason being that having predictive power about the future, could be completely caused by the partitioning.
-In [21] the correlation measured considered all possible states, and the measures were directly related to a causal effect.
-In addition, in [21] the shared information between the system with a node shifted over time ($I(S^{\tau}:s_{i}^{\tau+t})$) was considered.
+In [vanElteren2022] the correlation measured considered all possible states, and the measures were directly related to a causal effect.
+In addition, in [vanElteren2022] the shared information between the system with a node shifted over time ($I(S^{\tau}:s_{i}^{\tau+t})$) was considered.
 Applying this approach under a state partition $I(S^{\tau}:s_{i}^{\tau+t}|\langle S\rangle)$causes a violation of the data processing as information may flow from a node at a particular $t=t_{1}$ and then flow back to the node at $t=t2,t_{2}>t_{1}$.
@@
 For high $\beta$ values metastable behavior occurs in the kinetic Ising model.
-The temperature was chosen such that the statistical complexity [44] was maximized.
+The temperature was chosen such that the statistical complexity [Lopez-Ruiz1995a] was maximized.
 The statistical complexity $C$ is computed as
@@
 A.6.
-The noise parameter $\beta$ is set such that it maximizes the statistical complexity using numerical optimization (COBYLA method in scipy’s optimize.minimize module) [45].
+The noise parameter $\beta$ is set such that it maximizes the statistical complexity using numerical optimization (COBYLA method in scipy’s optimize.minimize module) [Virtanen2020].
 Figure A.6: (a) Statistical complexity ($C$), normalized system entropy ($H(S)$) and disequilibrium ($D(S)$) as a function of the temperature ($T=\frac{1}{\beta}$) for Krackhardt kite graph.
@@
 Let the minority state be 1 and the majority state be 0, the expectation of $s_{i}$ flipping from the majority state to the minority state is given as:
-$E[p(s_{i}=1|N)]_{p(N)}=\sum_{N_{i}\in N}p(N_{i})p(s_{i}=1|N_{i})\\ =\sum_{N_{i}\in N}\prod_{j}^{|N_{i}|}p(n_{j})p(s_{i}=1|N_{i})\\ =\sum_{N_{i}\in N}{n\choose k}f^{k}(1-f)^{n-k}p(s_{i}=1|f),$ (13)
+{dmath}
+[compact=-1000] E[ p(s_i = 1 | N ) ]_p(N) = ∑_N_i ∈N p(N_i) p(s_i = 1 | N_i)
+= ∑_N_i ∈N ∏_j^|N_i| p(n_j) p(s_i = 1 |N_i)
+= ∑_N_i ∈N (nk) f^k (1 - f)^n-k p(s_i = 1 | f),
 where $f$ is the fraction of nodes in the majority states, $n$ is the number of neighbors, $k$ is the number of nodes in state 0.